(Newtons laws/gravitation force/normal force/friction force, 4/ea.) A 60-N force pushes two objects next to each other with masses m1 8 kg, and m2 4 kg, across a horizontal solid surface. As shown below, the force is directed 30 below the horizontal. The coefficient of kinetic friction is 0.15, same for both objects. (a) What is the normal force on mu? Ans. (b) What is the normal force on m2? Ansi (c) What is the acceleration of the masses? Ans. (d) What is the force the m1 exerts on m2? Ans. (e) What is the force the m2 exerts on m1? Ans.
Answer #1

Free body diagram of m1 and m2

Fcos300 N1 Fsin300 f2 fı m1g

Applying Newton's law of motion along vertical direction for mass m1

N1= m1g + Fsin300-----------(1)

f1 = 0.15*N1

Applying Newton's law of motion along horizontal direction for mass m1

Fcos300 - N - f1 =  m1a

=> Fcos300 - N - 0.15*(m1g + Fsin300) =  m1a-----------(2)

Applying Newton's law of motion along vertical direction for mass m2

N2= m2g -----------(3)

f2= 0.15*N2

Applying Newton's law of motion along horizontal direction for mass m2

N - f2 =  m2a

=>N - 0.15*(m2g ) =  m2a-----------(4)

From equation 2 and 4

a = [Fcos300 - 0.15*(m1g +m2g + Fsin300)]/[m1 +m2 ]

From equation 4.

N = 0.15*(m2g ) + m2a

So answers are

(a) N1= m1g + Fsin300

(b) N2= m2g

(c) a = [Fcos300 - 0.15*(m1g +m2g + Fsin300)]/[m1 +m2 ]

(d) N = 0.15*(m2g ) + m2a

(e) N = 0.15*(m2g ) + m2a